OpenStudy (anonymous):

a runner went from 6m/s to 2m/s in 2 seconds what is the acceleration

OpenStudy (theeric):

Hi! Do you know what acceleration is?

OpenStudy (theeric):

It is the change in velocity per some change in time. So, what is your change in velocity? It is your final velocity minus your initial velocity. What is your change in time? It is directly given to you in the question. How do you say, "per some change in time"? You divide by that time. This is the definition of acceleration. I hope it helps!

OpenStudy (anonymous):

the acceleration is change in the velocity divided by change in time i.e \[a=\Delta v/\Delta t\] here a= 6-2/2 therefore a=4/2=2 m/s sqr